OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

If you have an equation of the line in the form y=ax+b then your slope will be a. Now, the line on the hipotesis is y=-3x-4, then your slope is (-3) If two lines are parallel, then their slopes are the same. Ergo, the line you're searching has slope (-3) Finally, you can find the equation of a line if you have the slope and a point. slope is "m", the point is (x1,x2) y-y1=m(x-x1) Then, you'll have y-6=-3(x-2) y=-3x+12